OpenStudy (anonymous):

What is 5 1/3 · (–4 4/5 )? A. –25 3/5 B. –20 4/15 C. 20 4/15 D. 25 3/5

OpenStudy (bibby):

ah we meet again panda... I'm going to assume you remember the process of multiplying mixed fractions: how do we get this into an improper fraction?:\[-4 \frac{ 4 }{ 5 }\]

OpenStudy (shamil98):

hint: (-4 x 5 ) -4 = -4 4 /5

OpenStudy (shamil98):

(-4 x 5) - 4 over 5 *

OpenStudy (anonymous):

\[-4 \frac{ 4 }{ 5 }= \frac{ 5·-4 }{ 5 }=\frac{ -20+4 }{ 5 }\]

OpenStudy (shamil98):

-20 - 4 = -24/5 -24/5 = -4 4/5

OpenStudy (bibby):

This time around, it's a bit different,. as one of the mixed numbers are negative. \[-4\frac{ 4 }{ 5 } = - (4 + \frac{ 4 }{ 5 }) = -(\frac{ 20 }{ 5 })\]

OpenStudy (bibby):

whoops the last bit didn't go through \[-(\frac{ 20 }{ 5 } + \frac{ 4 }{ 5 })\]

OpenStudy (anonymous):

so it would be \[-\frac{ 16 }{ 5 }\]

OpenStudy (bibby):

If you're following PEDMAS or something like that, you would do the bit in the parentheses first

OpenStudy (anonymous):

Thank you. :)

OpenStudy (bibby):

Glad to help :D Do you want to continue through to the end?

OpenStudy (anonymous):

yes

OpenStudy (bibby):

Les do dis. Since we know that -4 4/5 = -24/5 we have to find 5 1/3 as an improper fraction before we can multiply

OpenStudy (bibby):

Are you having any hesitations about getting 5 1/3 into fractional form?

OpenStudy (anonymous):

no i just had to get my food i am almost done

OpenStudy (bibby):

oh no problem. Eat and do whatever. I'll be here

OpenStudy (anonymous):

\[\frac{ 3*5 }{ 3 }=\frac{ 15+1 }{3}\]

OpenStudy (bibby):

yep. so now we have \[\frac{ 16 }{ 3 }* -\frac{ 24 }{ 5 }\]

OpenStudy (anonymous):

\[-\frac{ 384 }{ 24 }\]

OpenStudy (anonymous):

right?

OpenStudy (bibby):

yeah. and you can simplify that by dividing by 24 to get 16/1 but that's not especially necessary methinks

OpenStudy (anonymous):

so what would the answer be A B C D

OpenStudy (bibby):

whoops, you weren't right in the multiplication step. the denominator is 5*3 = 15 so it's -384/15

OpenStudy (anonymous):

so would it be A. \[-25 \frac{ 3 }{ 5 }\]

OpenStudy (bibby):

Yeah. Sorry for not crosschecking it earlier :/

OpenStudy (anonymous):

thats ok